The air pressure in my tires is about 35 psi when the temperature is 25 oC outside. What will the pressure in my tires be on a cold day with a temperature of -12 oC?


1
Expert's answer
2020-04-18T07:08:13-0400

As pV=nRT so volume and number of moles is constant so "\frac{p}{T}=constant"

"\dfrac{p_1}{T_1}=\dfrac{p_2}{T_2}\\ \dfrac {35}{273+25}=\dfrac{p_2}{273-12}\\p_2=30.65\ psi"